OpenStudy (yttrium):

If you flip a fair coin 1000 times, what is the expected value of the product of the number of heads and the number of tails?

OpenStudy (anonymous):

need help?

OpenStudy (anonymous):

okay buddy

OpenStudy (anonymous):

60%

hartnn (hartnn):

2 equally likely possibilities so expected number will also be same 500 and 500 25 0000

OpenStudy (yttrium):

no sir you got it wrong

hartnn (hartnn):

fair coin ... P(head) = P(tail) = 0.5 Expected Heads E(Heads) = P(Head)*1000 = 500 right till here ?

hartnn (hartnn):

ohh! you're asking E(Head*Tail) but I answered E(Head)*E(Tail) isn't it ?

hartnn (hartnn):

can't imagine P(head*tail) in one coin toss :P

hartnn (hartnn):

so if i am wrong I mis-interpreted the question! :O

hartnn (hartnn):

whats the correct approach ?

OpenStudy (anonymous):

Let \(X\) denote the number of heads, and \(Y\) the number of tails. Hence \(Y=1-X\). We want to find \(E[X(1-X)]\) @hartnn This is the same as \(E(X)-E(X^2)=E(X)-\bigg(V(X)+[E(X)]^2\bigg)\). Both \(X\) and \(Y\) are binomially distributed with \(p=0.5\) and \(n=1000\), and the expected value for a binomial distribution such as this is \(np\), or \(500\). The variance \(V(X)\) is given by \(np(1-p)\), or \(250\). \[E[X(1-X)]=500-250-500^2=-249,750\] Is this right @Yttrium ?

hartnn (hartnn):

what does negative Expectation Indicate ?

OpenStudy (anonymous):

Not really sure, I never had a good notion of what the product of random variables might represent unless they explicitly refer to things like length and width.